# Model Alignment Search

### v5

@@ 2.3 MAS Formulation
MAS can be thought of as an efficient form of model stitching (lenc2015modelstitching, ; bansal2021modelstitching, ) or a multi-model extension of DAS geiger2021 ; geiger2023boundless . In the former case, both stitching and MAS learn a mapping from the representations of one trained model to another. In the latter case, both DAS and MAS attempt to find a causally relevant subspaces in the representations that align with the variables from Directed Acyclic Graphs (DAGs). MAS further measures the degree to which the models’ aligned subspaces align with each other. All analyses are performed on trained NNs with $>99\%$ task accuracy with frozen weights. We note, however, that MAS can theoretically be used to align models that are not fully trained.
DAS tests the assumption that the hidden state, $h^{(i)}\in R^{d_{i}}$, for a single model $\psi_{i}$, can be written as an orthogonal rotation $h^{(i)}=Q_{i}z^{(i)}$, where $Q_{i}\in R^{d_{i}\times d_{i}}$ is a learned orthonormal matrix, $z^{(i)}\in R^{d_{i}}$ consists of contiguous subspaces that encode high-level variables from DAGs, and $d_{i}$ is the size of the hidden state.
$h^{(i)}=Q_{i}z^{(i)}=Q_{i}\begin{bmatrix}\vec{z}_{\text{var}_{1}}\\ \vec{z}_{\text{var}_{2}}\\ ...\\ \vec{z}_{\text{var}_{n}}\end{bmatrix}$ (3)
Where each $\vec{z}_{\text{var}_{k}}\in R^{d_{\text{var}_{k}}}$ is a column vector of length $d_{\text{var}_{k}}$ satisfying the relation $\sum_{k=1}^{n}d_{\text{var}_{k}}=d_{i}$. The benefit of this alignment is that it allows us understand $\psi_{i}$’s neural activity in terms of interpretable variables by allowing us to isolate and causally manipulate the value of each $\text{var}_{k}$ within the NN’s latent representations. MAS builds on DAS by measuring the degree to which $\psi_{1}$ and $\psi_{2}$ share the same $\vec{z}_{\text{var}_{k}}$ values for a given $\text{var}_{k}$. With this formulation, we can freely isolate and manipulate $\vec{z}_{\text{var}_{k}}$ encoded in $h^{(i)}$ using a causal intervention:
$h^{(i)}_{v}=Q_{i}^{-1}((1-D_{i,\text{var}_{k}})Q_{i}h^{(i)}_{trg}+D_{j,\text{var}_{k}}Q_{j}h^{(j)}_{src})$ (4)
@@
Where $i$ and $j$ can take on any model index in the set of all models considered, $Q_{i}$ is a scaled orthogonal rotation matrix for $\psi_{i}$, $D_{i,\text{var}_{k}}\in R^{d_{i}\times d_{i}}$ is a diagonal, binary matrix with $d_{\text{var}_{k}}$ non-zero elements used to isolate the dimensions corresponding to $\vec{z}_{\text{var}_{k}}$, $h^{(j)}_{src}$ is the source vector from which the subspace is harvested, $h^{(i)}_{trg}$ is the target vector into which activity is substituted, and $h^{(i)}_{v}$ is the resulting intervened vector that then replaces $h^{(i)}_{trg}$ in the target $\psi_{i}$’s processing, allowing the model to make causally intervened predictions. See Figure 1 for a visualization. In this work, we train $Q_{i}$ as the product $s_{i}U_{i}$ where $s_{i}$ is a learned scalar and $U_{i}$ is an orthonormal matrix.
Model stitching can be unified with DAS/MAS by considering a DAG where we assume that all behaviorally relevant information is encoded as a single variable, $\vec{z}_{full}$, and all extraneous, irrelevant information is encoded in a subspace $\vec{z}_{extra}$. In this case, $z^{(i)}=\begin{bmatrix}\vec{z}_{full}\\ \vec{z}_{extra}\end{bmatrix}$, and we can freely isolate and intervene upon $\vec{z}_{full}$ using Equation 4 once we have learned each $Q_{i}$ for each $D_{i,full}$ for each $\psi_{i}$.
@@ 2.4 MAS Training
MAS relies on the notion of counterfactual behavior to create intervention data to train and evaluate each $Q_{i}$. For a given DAG, we know what the DAG’s behavior will be after changing the value of one of its variables. The counterfactual behavior of the DAG is the resulting behavior of the DAG after changing a specific variable and keeping all other variables and aspects of the task unchanged. This counterfactual behavior can be used as a training signal for $Q_{i}$ using a standard Next-Token Prediction (NTP) autoregressive loss. $Q_{i}$ can equivalently learn any permutation of the subspaces in $z_{i}$, thus we can restrict our trainings to diagonal matrices $D_{i,\text{var}_{k}}$ that have contiguous, non-zero entries. It is then possible to treat $d_{\text{var}_{k}}$ as a hyperparameter in independent trainings, selecting the $Q_{i}$, $D_{i,\text{var}_{k}}$ pair with the best evaluation. Unless otherwise stated, we use values of $d_{\text{var}}$ from the better performing of 16 or 32, and we perform our causal interventions on individual time steps in the token sequences. We run the target model $\psi_{i}$ up a sampled timestep $t$ on the target sequence, using its latent representation at that point as the target vector, $h^{(i)}_{t,trg}$. We do the same on the source model $\psi_{j}$ to obtain the source vector, $h^{(j)}_{u,src}$, at timestep $u$ from a separate source sequence. We then construct $h^{(i)}_{t,v}$ using Equation 4. $\psi_{i}$ then continues its predictions starting from time $t$, using $h^{(i)}_{t,\text{v}}$ in place of $h^{(i)}_{t,trg}$. We perform batch gradient descent on $Q_{1}$ and $Q_{2}$ using the loss as follows for a single counterfactual sequence, $k$, of length $S_{k}$ using $i$ as the target model index and $p_{\theta_{i}}(x_{s})$ as the probability generated by $\psi_{i}$ for the counterfactual token label at step $s$ with N samples in the batch:
$\mathcal{L}_{i}^{(k)}(Q_{i},Q_{j})=-\frac{1}{S_{k}}\sum_{s=t}^{S_{k}}\log p_{\theta_{i}}\bigl{(}x^{(k)}_{s}\mid x^{(k)}_{<s},h^{(k),(i)}_{t,\text{v}}\bigr{)}$ (5)
$\mathcal{L}_{tot}=\sum_{i=1}^{2}\sum_{j=1}^{2}\frac{1}{N}\sum_{k=1}^{N}\mathcal{L}^{(k)}_{i}(Q_{i},Q_{j})$ (6)
For the LSTM architecture, we perform MAS on a concatenation of the $h$ and $c$ recurrent state vectors (hochreiter1997lstm, ). In the GRUs, we operate on the recurrent hidden state. In the transformers, we operate on the residual stream following the first transformer layer (referred to as the Layer 1 Hidden States in Supplementary Figure 4). Unless otherwise stated, we uniformly sample intervention time points, $t$ and $u$, from sequence positions containing demo tokens or response tokens (excluding B, T, and E tokens). Unless otherwise stated, we orthogonalize the rotation matrix using PyTorch’s default orthogonalization that uses the exponential of a skew symmetric matrix. See Appendix A.2 for more detail.
@@
MAS Evaluation: We can evaluate the quality of $Q_{1}$ and $Q_{2}$ using the accuracy of each model’s predictions on counterfactual outputs from held out intervention data. A trial is correct when all deterministic tokens are predicted correctly using the argmax over logits. Unless otherwise stated, we report the proportion of trials correct for the worst performing causal intervention pairing, ($i$, $j$), as the Interchange Intervention Accuracy (IIA). We calculate error bars as Standard Error over unique model seed pairings using 2 model seeds for each training variant. We include two additional seeds for within training type comparisons.
@@ 2.5 Mas Variants
Stepwise MAS: For all input embedding analyses we include a MAS variant that applies Equation 4 to multiple, contiguous time-steps in the target and source sequences from step 0 to $u$—contrasted to a single time step pair ($t$,$u$). This allows for meaningful transfer of information in anti-Markovian solutions (grant2024das, ). There is a question of what tokens to use to create each $h^{(i)}_{t,trg}$ before transfer. We show results where each target token comes from the original target sequence padded by the R token when $u$ exceeds the target sequence length. See Figure 1 for a visualization.
Counterfactual Latent MAS (CLMAS): In some settings we wish to examine systems for which we have neural recordings but no causal access (as is often the case in BNNs). We refer to causally inaccessible models as $\psi_{nacs}$ and accessible models as $\psi_{acs}$. In these cases, we can only train on the behavior of $\psi_{acs}$ but we can read the representations produced by $\psi_{nacs}$. In these cases, we can adapt Equation 6 to ignore the behavior of $\psi_{nacs}$ like so:
$\mathcal{L}_{acs}=\frac{1}{N}\big{(}\sum_{k=1}^{N}\mathcal{L}^{(k)}_{acs}(Q_{acs},Q_{nacs})+\sum_{k=1}^{N}\mathcal{L}^{(k)}_{acs}(Q_{acs},Q_{acs})\big{)}$ (7)
We will show that if we merely adapt MAS to train on $\psi_{acs}$’s behavior, the resulting $Q_{acs}$ and $Q_{nacs}$ do not achieve high IIA when performing causal interventions on $\psi_{nacs}$. To address this, we introduce an auxiliary loss function for the one-directional trainings. The auxiliary objective relies on Counterfactual Latent (CL) vectors, which we define as vectors that encode the causal variables that we would expect to exist in the intervened vector, $h^{(nacs)}_{trg,t}$ from Equation 4, following a causal intervention. We can obtain CL vectors by searching through a prerecorded $h^{nacs}$ dataset for situations and behaviors that are consistent with the information we wish to be encoded from the causal intervention. For example, if we have a DAG with variables $\text{v}ar_{y}$, $\text{v}ar_{w}$, and $\text{v}ar_{extra}$, and following a causal intervention we expect $h^{(nacs)}_{t,\text{v}}$ to have a value of $y$ for variable $\text{v}ar_{y}$ and $w$ for variable $\text{v}ar_{w}$, then the CL vector can be obtained from a pre-recorded representation $h^{(nacs)}_{CL}$ that has the same expected variable values: $\text{v}ar_{y}=y$ and $\text{v}ar_{w}=w$. The auxiliary loss $\mathcal{X}^{(k)}$ for a single intervention sample $k$ is composed of an L2 loss and a cosine loss using CL vectors as the ground truth:
$\mathcal{X}^{(k)}_{L2}$ $=$ $\frac{1}{2}||h_{t,\text{v}}^{(k),(nacs)}-h^{(k),(nacs)}_{CL}||^{2}_{2}$ (8)
$\mathcal{X}^{(k)}_{cos}$ $=$ $-\frac{1}{2}\frac{h_{t,\text{v}}^{(k),(nacs)}\cdot h^{(k),(nacs)}_{CL}}{||h_{t,\text{v}}^{(k),(nacs)}||_{2}\,||h^{(k),(nacs)}_{CL}||_{2}}$ (9)
@@
where $h^{(nacs)}_{t,\text{v}}$ is the intervened target vector for the causally inaccessible model, and $k$ denotes the index of the sample within the batch. The total CLMAS training loss is a weighted sum of the loss from Equation 7 and the auxiliary loss where $\lambda$ is a hyperparameter: $\mathcal{L}_{CMAS}=\lambda(\mathcal{X}_{L2}+\mathcal{X}_{cos})+(1-\lambda)\mathcal{L}_{acs}$
@@
Transferring Specific Variables Through our choice of counterfactual training behavior, we can explore alignment of specific types of representations between models. We use representations of number as a case study for their precise but general nature. To do this, we construct counterfactual sequences based on the assumption that the numeric information in $h_{src}$ is transferred to $h_{trg}$ and all other information is left unaffected. We assume that the models encode a single numeric variable in the Multi-Object, Same-Object, Modulo , and Rounding tasks grant2024das . In the arithmetic task, we assume there is a numeric representation for the number of remaining operations, Rem Ops, and another representation for the cumulative value, Cumu Val. See expanded details in Appendix A.7.
@@
Model Stitching: As baselines, we include causal intervention methods that we refer to as Latent Stitch and Stitch. Latent Stitch consists of a single orthogonal matrix $Q$ that learns to map $h_{u,src}^{(k),(acs)}$ to $h_{CL}^{(k),(nacs)}$ by minimizing the CL auxiliary loss from Equation 8 without the Equation 7 objective. Stitch consists of a single, possibly low rank, orthogonal matrix $Q$ trained in a single behavioral direction from $\psi_{nacs}$ to $\psi_{acs}$ without the CL auxiliary loss. Unless otherwise stated assume $Q$ is full rank and the IIA is reported from validation interventions in the trained direction, from $\psi_{nacs}$ to $\psi_{acs}$.
2.6 Gram-Schmidt Alignment Matrices
To save compute for analyses on large vectors, we present an intervention method that only learns the components specific to the subspace of interest. If we assume the number of dimensions in the aligned space is $n$, we can use the Gram-Schmidt process to orthogonalize and train only $n$ components of an orthogonal matrix for $\psi_{i}$. To do this, $n$ column vectors $\text{v}_{k}\in R^{d_{i}}$ are sampled from a normal distribution where $d_{i}$ is the latent dimensionality of $\psi_{i}$. We can then make an orthogonal set of vectors $\text{u}_{k}\in R^{d_{i}}$, and thus a partial orthogonal matrix $Q\in R^{m\times n}$, by subtracting out the projections of all preceding components from each $v_{i}$ where $\text{u}_{1}=\text{v}_{1}$:
$\text{proj}_{u}(\text{v})$ $=$ $\frac{\text{u}^{\top}\text{v}}{\text{u}^{\top}\text{u}}\text{u}$ (10)
$\text{u}_{k}$ $=$ $\text{v}_{k}-\sum_{i=1}^{k-1}\text{proj}_{u_{i}}(\text{v}_{k})$ (11)
$Q$ $=$ $\begin{bmatrix}\frac{u_{1}}{||u_{1}||}&\frac{u_{2}}{||u_{2}||}&...&\frac{u_{n}}{||u_{n}||}\end{bmatrix}$ (12)
Using partial rotation matrices $Q_{i}$ and $Q_{j}$ for $\psi_{i}$ and $\psi_{j}$, we can formulate a MAS (or DAS) intervention as follows (see Appendix A.8 for proof of Equation 13):
$h^{(i)}_{\text{v},t}=h^{(i)}_{trg,t}+Q_{i}(-Q_{i}^{\top}h^{(i)}_{trg,t}+Q_{j}^{\top}h^{(j)}_{src,u})$ (13)

### v6

@@ 3.3 MAS Formulation
At a high level, MAS can be thought of as a combination of model stitching and DAS. MAS causally measures the degree to which multiple models’ behavioral subspaces align with each other.
Using our notation from Sections 2.1 and  2.2, MAS learns an invertible alignment function, $\mathcal{A}_{\psi_{i}}:\mathbb{R}^{d_{\psi_{i}}}\rightarrow\mathbb{R}^{d_{\psi_{i}}}$ for each $\psi_{i}$ of $\mathbb{N}$ models in the analysis, such that $\mathcal{A}_{\psi_{i}}(h_{\psi_{i}})=z_{\psi_{i}}$, and within each $z_{\psi_{i}}$, MAS performs interchange interventions to examine the degree to which the subspaces are causally interchangeable for all included models. It does this using both between and within model interchange interventions. Effectively, MAS asks, does $\vec{z}_{\psi_{i},\text{var}_{k}}=\vec{z}_{\psi_{j},\text{var}_{k}}\forall\,(i\in\mathbb{N},j\in\mathbb{N})$? We limit our empirical analyses to $\mathbb{N}=2$ and we only examine cases where $\mathcal{A}_{\psi_{i}}=Q_{\psi_{i}}=a_{\psi_{i}}U_{\psi_{i}}$ where $a_{\psi_{i}}$ is a scalar and $U_{\psi_{i}}\in R^{d_{\psi_{i}}\times d_{\psi_{i}}}$ is an orthogonal matrix.
With these intutions, we can generalize Equation 3 to the multi-model case for an intervention on the subspace for $\text{var}_{k}$:
$h_{\psi_{i}}^{(v)}=Q_{\psi_{i}}^{-1}((1-D_{\psi_{i},\text{var}_{k}})Q_{\psi_{i}}h^{(trg)}_{\psi_{i}}+D_{\psi_{j},\text{var}_{k}}Q_{\psi_{j}}h^{(src)}_{\psi_{j}})$ (6)
@@
Where $i$ and $j$ can be any model index in the set of all models considered, $Q_{\psi_{i}}$ is a scaled orthogonal rotation matrix for $\psi_{i}$, $D_{\psi_{i},\text{var}_{k}}\in R^{d_{\psi_{i}}\times d_{\psi_{i}}}$ is a diagonal, binary matrix with $d_{\text{var}_{k}}$ non-zero elements used to isolate the dimensions corresponding to $\vec{z}_{\text{var}_{k}}$, $h^{(src)}_{\psi_{j}}$ is the source vector from which the subspace is harvested, $h^{(trg)}_{\psi_{i}}$ is the target vector into which activity is substituted, and $h^{(v)}_{\psi_{i}}$ is the resulting intervened vector that replaces $h^{(trg)}_{\psi_{i}}$ in $\psi^{(trg)}_{i}$’s processing, allowing the model to make causally intervened predictions. See Figure 1(a) for a visualization.
For many of our analyses, we will perform MAS without segregating the behaviorally relevant subspace. Concretely, this can be formalized as a causal abstraction that uses a single Full variable in which all behaviorally relevant information is encoded, $\vec{z}_{full}$, and all extraneous information is encoded in $\vec{z}_{extra}$, the behavioral null-space and dormant subspaces. For completion: $z_{\psi_{i}}=\begin{bmatrix}\vec{z}_{\psi_{i},full}\\ \vec{z}_{\psi_{i},extra}\end{bmatrix}$, and we will freely isolate and manipulate $\vec{z}_{\psi_{i},full}$ using Equation 3.
@@ 3.4 MAS Training
The MAS training procedure is similar to DAS in that the training loss is created using the output distribution from $\psi_{i}^{(trg)}$ conditioned on $h^{(v)}_{\psi_{i}}$ as predictions and the counterfactual behavior of the causal abstraction as labels. In the case of $\vec{z}_{full}$, the counterfactual behavior is the same as that of model stitching from Equation 1, where the training labels are the behavior from $\psi_{src}$ following the intervention. Using $y$ to denote the counterfactual behavior labels in the loss for a single intervention, the loss for a single intervention direction is as follows:
$\mathcal{L}(Q^{(trg)}_{\psi_{i}},Q^{(src)}_{\psi_{j}})\;=\;\mathbb{E}_{x,y\sim\mathcal{D}}\big{[}\ell\big{(}f_{\psi_{i}}(h^{(v)}_{\psi_{i}})\,\|\,y\big{)}\big{]}$ (7)
In Sections 3.5 and  4.5 we will address cases in which we only include a subset of the intervention direction permutations in the training loss. However, unless otherwise stated, we default to averaging over all permutations. The total loss for a single matrix $Q_{\psi_{i}}$ is as follows:
$\mathcal{L}_{total}(Q_{\psi_{i}})\;=\;\frac{1}{2\mathbb{N}-1}\sum_{j=1}^{\mathbb{N}}\big{(}\mathcal{L}(Q_{\psi_{i}},Q_{\psi_{j}})+\mathbf{1}_{\{i\neq j\}}\mathcal{L}(Q_{\psi_{j}},Q_{\psi_{i}})\big{)}$ (8)
When using auto-differentiation frameworks (Paszke et al., 2019), we simply use the gradient from the mean of the losses in all directions:
$\mathcal{L}_{total}\;=\;\frac{1}{\mathbb{N}^{2}}\sum_{i=1}^{\mathbb{N}}\sum_{j=1}^{\mathbb{N}}\mathcal{L}(Q_{\psi_{i}},Q_{\psi_{j}})$ (9)
It is crucial to include interventions where $i=j$ as this adds a soft constraint to the alignment that encourages the separation of $\vec{z}_{full}$ from $\vec{z}_{extra}$.
@@
MAS Evaluation: We can evaluate the quality of each $Q_{\psi_{i}}$ using the accuracy of the models’ predictions on held out counterfactual data following the interventions. We always report the IIA of the worst performing causal direction in the analysis. In the numeric tasks experiments, a trial is considered correct when all deterministic tokens are predicted correctly using the argmax over logits. For the LLM toxicity experiments, we report token prediction accuracy rather than trial accuracy. We calculate error bars as Standard Error over unique model seed pairings. We include two additional seeds for within training type comparisons in the numeric task experiments, and three seeds for each of the DeepSeek finetunings.
@@ 3.5 Mas Variants
We explore a number of variations on the MAS procedure as a means of demonstrating how to use MAS in different situations for different purposes. We enumerate these variants in this section.
Stepwise MAS: Up to this point, we have described a form of MAS that performs interventions on latent vectors at individual time points in the sequence. We include an exploration of a variant in which Equation 6 is applied at each time point from the beginning of the sequence up to some time point $t$. For all embedding layer analyses, we use this variant. See Figure 1(b) for a visual depiction.
Counterfactual Latent MAS (CLMAS): An interesting use case for neural similarity techniques is the comparison of ANNs with BNNs. An issue with causal comparisons of this nature is that we often have no causal access to the BNN in the analysis. We do, however, usually have neural recordings from the BNN. Using a tilde to denote causally inaccessible models, $\tilde{\psi}_{i}$, we can simulate such ANN-BNN comparisons by omitting causal interventions using $\tilde{\psi}_{i}$ as the target model from the training loss. We will assume, however, that we can still read from the latent vectors produced by $\tilde{\psi}_{i}$—analogous to a neural recording. We adapt Equation 8 for $\psi_{k}$, where $k\neq i$, to ignore the behavior of $\tilde{\psi}_{i}$ as follows:
$\mathcal{L}_{accessible}(Q_{\psi_{k}})=\frac{1}{2\mathbb{N}-1}\big{(}\mathcal{L}(Q^{(trg)}_{\psi_{k}},Q^{(src)}_{\tilde{\psi}_{i}})+\sum_{j=1}^{\mathbb{N}}\mathbf{1}_{\{j\neq i\}}\big{(}\mathcal{L}(Q^{(trg)}_{\psi_{k}},Q^{(src)}_{\psi_{j}})+\mathbf{1}_{\{j\neq k\}}\mathcal{L}(Q^{(trg)}_{\psi_{j}},Q^{(src)}_{\psi_{k}})\big{)}\big{)}$ (10)
And the loss for $\tilde{\psi}_{i}$ is $\mathcal{L}_{inaccessible}(Q_{\tilde{\psi}_{i}})=\frac{1}{2\mathbb{N}-1}\sum_{j=1}^{\mathbb{N}}\mathbf{1}_{\{j\neq i\}}\mathcal{L}(Q^{(trg)}_{\psi_{j}},Q^{(src)}_{\tilde{\psi}_{i}})$.
[…]
As an example, if we have a causal abstraction with variables $\text{v}ar_{1}$, $\text{v}ar_{2}$, and $\text{v}ar_{extra}$, and, we expect $h^{(v)}_{\tilde{\psi}_{i}}$ to have a value of $y$ for variable $\text{v}ar_{1}$ and $w$ for variable $\text{v}ar_{2}$ after applying Equation 6, then a valid CL vector, $h^{(CL)}_{\tilde{\psi}_{i}}$, would be a pre-recorded representation in which the causal abstraction has labeled $h^{(CL)}_{\tilde{\psi}_{i}}$ to have the variable values: $\text{v}ar_{1}=y$ and $\text{v}ar_{2}=w$.
The auxiliary loss $\mathcal{X}$ for a single intervention sample is composed of an L2 loss and a cosine loss using CL vectors as the ground truth:
[…]
@@
where $h^{(v)}_{\tilde{\psi}_{i}}$ is the intervened target vector for the causally inaccessible model. The total CLMAS training loss is a weighted sum of the loss from Equation 10 and the auxiliary loss where $\epsilon$ is a hyperparameter: $\mathcal{L}_{CL}=\epsilon(\mathcal{X}_{L2}+\mathcal{X}_{cos})+(1-\epsilon)\mathcal{L}_{accessible}$. We show results from the best performing training of $\epsilon$ values 0.5, 0.89, 0.94, and the best $d_{full}$ out of 32, 64, and 128 dimensions.
@@
Transferring Specific Variables Through our choice of counterfactual training behavior, we can explore alignment of specific subspaces between models that do not possess the same domain or co-domain. We focus on representations of number as a case study for their precise but general nature. To do this, we narrow the MAS interchange interventions to $\vec{z}_{\psi_{1},numeric}$ and $\vec{z}_{\psi_{2},numeric}$ by constructing counterfactual sequences for each $\psi$’s co-domain. We assume that the models encode a single numeric variable in the Multi-Object, Same-Object, Modulo, and Rounding tasks Grant et al. (2025). In the arithmetic task, we assume there is a numeric representation for the number of remaining operations, Rem Ops, and another representation for the cumulative value, Cumu Val. See expanded details in Appendix B.7.
@@
Model Stitching: As baselines, we include causal intervention methods that we refer to as Latent Stitch and Stitch. Latent Stitch consists of a single orthogonal matrix $Q$ for two models, $\psi_{1}$ and $\psi_{2}$, that learns to map $h^{(src)}_{\psi_{1}}$ to $h^{(CL)}_{\psi_{2}}$ by minimizing the CL auxiliary loss from Equation 11 without including the behavioral objective Equation 10. Stitch consists of a single, possibly low rank, orthogonal matrix $Q$ trained in a single behavioral direction from $\tilde{\psi}_{1}$ to $\psi_{2}$ without using the CL auxiliary loss. Unless otherwise specified, assume $Q$ is full rank and the IIA is reported from validation intervention data in the $\tilde{\psi}_{1}$ to $\psi_{2}$ direction only.